# Slow magnetic quantum oscillations in the $c$-axis magnetoresistance of UTe2

### v1

@@ V Acknowledgments
We thank A. Eaton and M. Kartsovnik for fruitful discussions. We are grateful for the support by A. Mackenzie and the department of Physics of Quantum Materials at the MPI for Chemical Physics of Solids in Dresden, Germany. We acknowledge support by the Institute for Materials Research at Tohoku University, Japan. We acknowledge support from the Deutsche Forschungsgemeinschaft (DFG) Grant HE 8556/3-1 and the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter—ct.qmat (EXC2147, Project No. 390858490). We acknowledge the support of the HLD at HZDR, member of the European Magnetic Field Laboratory (EMFL).

### v2

@@ V Acknowledgments
We thank A. Eaton and M. Kartsovnik for fruitful discussions. We are grateful for the support by A. Mackenzie and the department of Physics of Quantum Materials at the MPI for Chemical Physics of Solids in Dresden, Germany. We acknowledge support by the Institute for Materials Research at Tohoku University, Japan. We acknowledge support from the French National Agency for Research ANR within the projects FRESCO (Project No. ANR-20-CE30-0020) and SCATE (Project No. ANR-22-CE30-0040). We acknowledge support from the Deutsche Forschungsgemeinschaft (DFG) Grant No. HE 8556/3-1 and the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter—ct.qmat (EXC2147, Project No. 390858490). We acknowledge the support of the HLD at HZDR, member of the European Magnetic Field Laboratory (EMFL).